Hydrolysis of a compound B(formula C2H6Cl12) was carried out by refluxing with excess potassium hydroxide solution.The resulting solution was acidified with dilute nitric acid and excess silver nitrate added to precipitate the chloride ions as silver chloride,0.718g of silver chloride was formed from 0.805g of B. (a)calculate the number of chlorine atoms present in each molecule of B which are liberated by hydrolysis. (b)suggest an explanation for the result obtained in (a). (c)give a possible structure


1
Expert's answer
2021-05-08T23:53:54-0400

"Mass = 0.805-0.718=0.087"


"Moles = \frac{0.087}{35.453}=2.45×10^{-3}"